### v1

# SplatVoxel: History-Aware Novel View Streaming without Temporal Training

Abstract
We study the problem of novel view streaming from sparse-view videos, which aims to generate a continuous sequence of high-quality, temporally consistent novel views as new input frames arrive. However, existing novel view synthesis methods struggle with temporal coherence and visual fidelity, leading to flickering and inconsistency. To address these challenges, we introduce history-awareness, leveraging previous frames to reconstruct the scene and improve quality and stability. We propose a hybrid splat-voxel feed-forward scene reconstruction approach that combines Gaussian Splatting [22] to propagate information over time, with a hierarchical voxel grid for temporal fusion. Gaussian primitives are efficiently warped over time using a motion graph that extends 2D tracking models to 3D motion, while a sparse voxel transformer integrates new temporal observations in an error-aware manner. Crucially, our method does not require training on multi-view video datasets, which are currently limited in size and diversity, and can be directly applied to sparse-view video streams in a history-aware manner at inference time. Our approach achieves state-of-the-art performance in both static and streaming scene reconstruction, effectively reducing temporal artifacts and visual artifacts while running at interactive rates (15 fps with 350ms delay) on a single H100 GPU.
footnotetext: * Work done during an internship at Google.
Figure 1: (Top) Per-frame reconstruction methods, which produce an independent scene reconstruction at each time step, are prone to flickering artifacts. (Bottom) In constrast, our history-aware novel view streaming model merges previous and current frame information, allowing us to better model occluded regions and improve temporal stability. Our method achieve state-of-the-art visual quality and temporal consistency, and runs in interactive rate (15 fps with a 350ms delay) on two view inputs of resolution $320\times 240$.
1 Introduction
Novel view streaming with sparse views is a challenging yet essential problem with applications in augmented and virtual reality (AR/VR). The goal is to reconstruct and render temporally consistent novel views when new frames continuously come in. While recent advances in novel view synthesis, such as 3D Gaussian Splatting (3DGS) [22], enable real-time rendering, they typically require minutes of per-scene optimization using densely sampled views to construct a scene representation. This computational overhead, combined with the need for dense camera coverage, makes them impractical for online streaming reconstruction with sparse camera setups, where each frame must be processed within a reasonable time budget for video output. Feed-forward reconstruction methods [14, 5, 79] trained with large-scale scene datasets offer a more efficient alternative by enabling sparse view reconstruction and rendering without per-scene optimization. However, these methods operate on each frame independently at inference time, omitting temporal information. As a result, they usually struggle with flickering artifacts and occlusion handling.
In this paper, we tackle the problem of novel view streaming with history-awareness, leveraging information from previous frames. This is particularly crucial in sparse-view settings, where per-frame data is often insufficient and occlusions are common. By incorporating history, we can enhance reconstruction quality and mitigate temporal flickering artifacts, significantly improving the user experience. However, working on multi-view dynamic scene reconstruction presents its own challenges. Large-scale, multi-view video datasets suitable for training are very limited, and efficiently modeling temporal correspondences is nontrivial. To this end, we present a novel feed-forward scene reconstruction method that effectively leverages historical information for streaming reconstruction without requiring training on the currently limited multi-view video datasets. Our approach introduces a hybrid Splat-Voxel representation that combines the advances of Gaussian Splatting [22] and an efficient sparse voxel transformer. Specifically, we adopt Gaussian Splatting as a moveable point-based representation, enabling real-time splatting-based rendering and facilitating deformation across frames. We warp the historical feed-forward reconstruction of Gaussian Splatting using an efficient motion uplifting algorithm, which leverages existing 2D motion priors from pretrained models [10, 64]. To integrate historical information carried from warped Gaussian primitives with new observations from the current frame, we introduce a sparse voxel transformer with a coarse-to-fine voxel representation, enabling effective information fusion within a temporally shared Euclidean space. Our method achieves state-of-the-art performance in terms of rendering quality and temporal consistency on several multi-view video datasets under the sparse camera setup, and can reach interactive rate (15fps with 350ms delay) under a single H100 GPU. In summary, the main contributions of our paper are as follows:
  • We propose a hybrid Splat-Voxel feed-forward reconstruction framework that leverages historical information to enable novel view streaming, without relying on multi-view video datasets for training.
  • We develop an efficient sparse voxel transformer with a coarse-to-fine voxel representation, outperforming existing feed-forward Gaussian splatting methods.
  • Experiment results demonstrate that our proposed framework enhances novel view synthesis for streaming scene reconstruction with better visual quality and reduced temporal artifacts through history-aware modeling.
2 Related Work
Novel View Synthesis
Image-based rendering approaches for novel view synthesis (NVS) predict the appearance of a scene from novel viewpoints given a set of input images. A key design decision is the choice of 3D scene representation to establish correspondences and capture structure. Popular representations include multi-plane images [13, 82, 4, 71, 65, 72, 27], neural fields [40, 56], voxel grids [61, 41, 50], and Gaussian splatting [22]. NVS methods can be broadly divided into those that optimize scene representation at test time and those that directly predict it through a feed-forward network [67, 59, 52, 78, 5, 12]. Recently, transformers have been used to reason about scenes without explicit 3D bias [20, 54, 53] or in combination with differentiable renderers in a hybrid approach [79, 17]. Our method follows the latter paradigm, leveraging a transformer backbone that integrates voxel grids and 3D Gaussians for scene representation while operating in a data-driven feed-forward manner to achieve streaming applications.
Reconstructing Dynamic Scenes
Handling dynamic motions in novel view synthesis (NVS) is challenging due to occlusions and the need for temporal consistency. Dynamic scene reconstruction is typically addressed in multi-view setups with fixed cameras [75, 57, 73, 8, 38] or monocular settings where both the camera and scene move [43, 28, 74, 29, 68]. Most methods optimize on the entire video sequences, relying on learned or hand-crafted priors such as optical flow [64, 36], tracking [10], and depth estimation [44, 76, 48, 49] for supervision [28, 74]. To regularize motion, these methods often incorporate basis functions [29, 68, 30] or motion priors [47], while others explicitly model static and dynamic components separately [21, 32, 35]. These techniques support applications such as animation [63, 18, 47], tracking [68, 39], and semantic understanding [33]. Some methods accelerate reconstruction by preserving past reconstructions for faster optimization [69, 25, 62], but they remain computationally expensive and impractical for real-world, sparse-view streaming scenarios. Our method uses pretrained models (specifically, point tracking) for motion priors. However, unlike typical methods that optimize on the entire video sequence, our model operates causally and online, using only previously received frames.
Online 3D Reconstruction
Feed-forward approaches achieve online 3D reconstruction by predicting novel views in a single pass without per-scene optimization. However, many rely solely on the current frame [14, 34, 51, 31], neglecting temporal information, which leads to flickering and occlusion artifacts. Some methods [51, 31] train on large-scale dynamic video datasets, often incorporating synthetic data due to the scarcity of real-world multi-view videos. However, they assume access to full sequences at inference, limiting their applicability in real-time settings. Our approach bridges these gaps by combining the efficiency of feed-forward inference with history-awareness. Unlike optimization-based methods, we do not require future frames, and unlike standard feed-forward models, we leverage past frames to improve temporal consistency. By generalizing from static scene training to dynamic video inference without additional temporal supervision, our method enables efficient, real-time novel view streaming from sparse multi-view inputs.
Table 1: Method Comparison. Unlike existing baselines, our framework uniquely supports the three key requirements of the novel view streaming task targeted in this paper: fast reconstruction with interactive rate , sparse view inputs, and history awareness. In particular, our streaming approach only uses past frames, in constrast to DeformableGS [77] and DeformableNerf [43] that optimize over $\textit{all}^{*}$ frames. The history-awareness in our model mitigates temporal flickering and occlusion artifacts, and is also able to support unlimited sequence lengths.
Method  Interactive Speed  Sparse-view Input  History Awareness
NeRF [40] $\times$ $\times$ $\times$
InstantNGP [41] $\times$ $\times$ $\times$
3DGS [22] $\times$ $\times$ $\times$
[…]
DeformableNerf [43] $\times$ $\checkmark^{*}$
[…]
GS-LRM [79] $\times$
Quark [14] $\times$
Ours
Figure 2: Method Overview. Our hybrid SplatVoxel model first extracts input image features using a multi-view transformer, which outputs pixel-aligned Gaussian splats for each input image with associated features. The splat features are then deposited onto a coarse-to-fine voxel grid using the decoded position, and a secondary sparse voxel transformer processes the grid features to output final Gaussian parameters. To merge history, we compute the triangulated scene flow from the input views and perform keypoint guided deformations. These deformed splats can be treated identically to the input-aligned splats, and be similarly deposited into voxel grid to merge the previous state with the current state. Our system is trained only on static scenes, and can generalize to dynamic scenes during inference.
3 Online Novel View Streaming
Given sparse (e.g., 2–4 views) RGB video streams from fixed and calibrated cameras, we aim to achieve online novel view synthesis, using only temporal information from past frames. The key objectives for our approach include efficient reconstruction (i.e. $>10$ fps interactive rate) and effective reuse of past information. History awareness is essential for our task given the sparse coverage of the input cameras, which can lead to artifacts such as temporal flickering and degraded reconstruction in occluded regions. Rather than processing each frame independently, using information from prior frames improves the per-frame reconstruction quality and ensures temporal consistency.
Challenges
Optimization-based methods [40, 2, 41, 22, 73] are not well-suited for this task, as they require extensive optimization over entire video sequences. Additionally, they perform poorly with sparse views and face challenges in representing long video sequences due to memory and capacity constraints. Recent feed-forward reconstruction methods [5, 6, 79, 14] have shown promise in efficient sparse-view reconstruction of static scene by utilizing large networks trained on large-scale multi-view datasets. However, their lack of explicit temporal modeling makes them prone to temporal flickering and occlusions in video inputs, resulting in suboptimal performance for our novel view streaming task.
Table 1 highlights the differences among these prior methods, none of which address novel view streaming or incorporate historical information. Integrating history-awareness into existing frameworks is particularly challenging. A major obstacle in training history-aware feed-forward networks is the limited availability of high-quality multi-view video datasets. Moreover, naive approaches, such as continuously warping historical scene reconstructions to later frames, fail to capture newly appearing objects and changes in appearance—not to mention that accurate 3D scene motion estimation from sparse multi-view RGB inputs remains challenging and prone to accumulated errors, ultimately degrading reconstruction fidelity.
Our Solution
To address these challenges, we propose a novel feed-forward reconstruction framework that generalizes to temporal video streams at inference time, despite being trained solely on multi-view static scenes. Our approach leverages a hybrid point-voxel representation, combining the flexibility of point-based methods with the structured spatial organization of voxels, drawing inspiration from hybrid Lagrangian-Eulerian methods [19] commonly used in physics simulation.
Our two-stage pipeline first constructs a movable point-based (Lagrangian) representation, which is then transferred into a voxel-based (Eulerian) representation. This design naturally extends static scene reconstruction to history-aware dynamic scene reconstruction, where historical information is retained within the Lagrangian component and continuously integrated with new data on the global Eulerian grid, following the principles of hybrid Lagrangian-Eulerian methods.
For our novel view streaming task, we employ Gaussian Splatting [22] as the Lagrangian component and a coarse-to-fine sparse voxel grid as the Eulerian component. A multi-view transformer learns to generate Lagrangian Gaussian primitives from multi-view RGB images, while a voxel transformer fuses information within the Eulerian grid in latent space. After training on a large-scale static scene dataset, our method enables online novel view streaming without additional temporal training. This is achieved by warping Gaussian primitives using an embedded deformation graph [60] derived from uplifted 2D tracks [10] and integrating the warped primitives with the current frame using the voxel transformer.
4 Method
In this section, we describe our SplatVoxel feed-forward scene reconstruction framework (Sec 4.1), and how it can generalize at inference time for zero-shot history-aware 4D novel view streaming (Sec 4.2).
4.1 Efficient SplatVoxel Transformer
Our two-stage feed-forward reconstruction model fuses information across 2D and 3D domains by first using a multi-view transformer to process 2D images into Gaussian splat primitives (we refer to a 3D Gaussian primitive as a “splat” for brevity here and in the following sections). We then deposit the splats on a voxel grid and train a sparse coarse-to-fine voxel transformer to refine the Gaussian primitives.
4.1.1 Multi-view Transformer
Given $N$ RGB images of resolution $H\times W$ with known camera parameters, we extract multi-view latent features and train a transformer following GS-LRM [79]. We concatenate the RGB value with per-pixel Plücker ray embeddings [45]. Next, we extract features from non-overlapping $p\times p$ image patches [11] using a single-layer CNN, resulting in patch features with channel dimension of $C$. The resulting multi-view image patch features are concatenated into a set of $N\times HW/p^{2}$ tokens, which we denote as $\{\mathbf{X}\}$. The multi-view (MV) transformer [66] then processes image tokens $\mathbf{X}\}$ to capture the underlying geometric and appearance, producing a set of latent features $\{\mathbf{Z}\}$:
$\centering\{\mathbf{Z}\}=\mathrm{Transformer}_{MV}(\{\mathbf{X}\})\@add@centering$ (1)
We unpatchify the processed latent features to Gaussian primitives with a linear layer applied to each latent:
$\centering\{\mathbf{G}_{k},\mathbf{F}_{k}\}=\mathrm{Linear}(\{\mathbf{Z}\})\@add@centering$ (2)
where $\mathbf{G}_{k}$ is a Gaussian primitive aligned with each input pixel ($k=1...,NHW$) which can be rendered using Gaussian splatting [22], and $\mathbf{F}_{k}$ is an associated feature vector for each Gaussian primitive.
4.1.2 Splat-to-Voxel Transfer
While previous feed-forward Gaussian Splatting methods [6, 79, 5] typically represent the scene by simply concatenating per-pixel splats predicted from different views, we propose a 3D-aware refinement step on splats $\mathbf{G}_{k}$ and associated features $\mathbf{F}_{k}$ by leveraging an intermediate voxel-based representation. This Splat-to-Voxel Transfer is achieved by first performing Splat Depositing onto a voxel grid, followed by Splat Fusion in Euclidean space to produce the input for our sparse voxel transformer.
Splat Depositing
Given voxel grid $V$, we deposit each Gaussian primitive $\mathbf{G}_{k}$ to the set of $\{\mathbf{M}_{k}\}$ of nearest voxels based on the position $\mu_{k}$ using a distance kernel $\mathcal{K}$. That is, for voxel $V_{j}$, its weight associated with $\mathbf{G}_{k}$ is:
$\centering w_{jk}=\begin{cases}\frac{\mathcal{K}(\mu_{k}-\mathbf{x}_{j})}{\sum_{j^{\prime}\in\{\mathbf{M}_{k}\}}\mathcal{K}(\mu_{k}-\mathbf{x}_{j^{\prime}})},&\text{if }j\in\{\mathbf{M}_{k}\},\\ 0&\text{otherwise}.\end{cases}\@add@centering$ (3)
where $\mathbf{x}_{j}$ represents the center of the $j$-th voxel (details in supplementary). We choose the adjacent 8 voxels for each splat in our implementaion. We note that the same splat depositing procedure can be used for Gaussian primitives originating from multi-view images, or warped from previous frames as will be detailed in Sec. 4.2. This seamlessly integrates all incoming splats which will be used to generate refined primitives with the downstream voxel transformer.
Splat Fusion
Each voxel ${V_{j}}$ then accumulates the deposited splat attributes using opacity $\alpha$ as weighting. The corresponding voxel attribute $\mathbf{E}_{j}$ at voxel $V_{j}$ is then the normalized weighted sum of the splat features $\mathbf{F}_{k}$ in the voxel:
$\mathbf{E}_{j}=\frac{1}{\mathbf{W}_{j}}\sum_{k\in\{\mathbf{G}_{k}\}}w_{jk}\alpha_{k}\mathbf{F}_{k},\quad\mathbf{W}_{j}=\sum_{k\in\{\mathbf{G}_{k}\}}w_{jk}\alpha_{k}$ (4)
where $w_{jk}$ is interpolation weight computed in Eqn. 3. We use the notation $k\in\{\mathbf{G}_{k}\}$ to denote an index into the set of all primitives, with associated weight, opacity, and feature.
Coarse-to-fine voxel hierarchy
Once the splat features are deposited onto the voxels, we form the coarse-to-fine voxel hierarchy. Using the previously computed voxel features $\mathbf{E}_{j}$ as the fine voxel features, we now compute coarse voxel features $\mathbf{E}^{c}_{i}$. Let $\mathbf{N}_{i}$ be the set of fine voxels associated $i$-th coarse voxel. We derive coarse voxel features using a shallow multi-layer perceptron (MLP) and set coarse voxel weights as the sum of the associated fine voxel weights:
$\mathbf{E}^{c}_{i}=\mathrm{MLP}_{coarse}\left([\mathbf{E}_{j}\mid j\in\mathbf{N}_{i}]\right),\quad\mathbf{W}_{i}=\sum\limits_{j\in\mathbf{N}_{i}}\mathbf{W}_{j}$ (5)
We leverage geometric cues from coarse voxel weights $\mathbf{W}_{i}$, which encapsulate the spatial distribution and opacity of Gaussian primitives within each voxel, to efficiently cull empty and insignificant regions of the scene. We sort the coarse-level voxels by their voxel weights and retain only the top 20% with the largest weights as a sparse voxel representation of the scene. This sparsification is crucial for the downstream voxel transformer, and reduces the input token count to fewer than $10K$, which allows the voxel transformer to run in $15$ ms.
4.1.3 Sparse Voxel Transformer
From coarse voxel features, we treat each voxel feature as a token and reshape to 1D vectors. We train a transformer to process the set of coarse voxel features $\mathbf{E}^{c}$:
$\centering\{\mathbf{O}^{c}\}=\mathrm{Transformer}_{voxel}(\{\mathbf{E}^{c}\})\@add@centering$ (6)
The processed latent features $\{\mathbf{O}^{c}\}$ are then replicated from the coarse voxel grids within their corresponding fine voxel grids, producing latent features $\{\mathbf{O}\}$ at the fine voxel resolution. A shallow MLP subsequently generates refined Gaussian primitives by integrating the initial fine-level voxel attributes $\mathbf{E}_{j}$ with the fine-level transformer latent $\mathbf{O}_{j}$.
$\centering\mathbf{G}^{\prime}_{j}=\mathrm{MLP}_{fine}([\mathbf{O}_{j},\mathbf{E}_{j}])\@add@centering$ (7)
where $\mathbf{G}^{\prime}_{j}\in\mathbb{R}^{l}$ represents the predicted Gaussian primitives around the sparsified fine-level voxel grids, with $l$ denoting the number of rendering parameters.
4.1.4 Training
We train the SplatVoxel feed-forward networks on multi-view static scene datasets [82, 37] using a photometric loss combining Mean Squared Error (MSE) and perceptual LPIPS [81]:
$\mathcal{L}=\mathcal{L}_{\text{MSE}}(\hat{I},I)+\lambda\mathcal{L}_{\text{LPIPS}}(\hat{I},I)$ (8)
where $\hat{I}$ denotes ground-truth target image and $I$ the corresponding rendered target image from the predicted Gaussian primitives, which come from either the multi-view transformer or voxel transformer. We first train the multi-view transformer, then jointly both multi-view and voxel transformers together. We set $\lambda$ to 0.5 for the multi-view transformer and 4.0 for the voxel transformer. Training details are provided in the supplementary.
Table 2: Re10K Reconstruction Metrics. We train our model on the RealEstate10K dataset and compare it with feed-forward Gaussian Splatting methods. Baseline metrics are obtained from their offical papers.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
pixelSplat [5] 26.09 0.863 0.136
MVSplat [6] 26.39 0.869 0.128
GS-LRM [79] 28.10 0.892 0.114
Ours 28.47 0.907 0.078
Figure 3: Static scene reconstruction. Qualitative comparison of our method with recent feed forward GS-LRM [79] on RealEstate10k [82]. Our method preserves high fidelity reconstructions and sharp details.
4.2 History-aware Streaming Reconstruction
In this section, we describe how our trained feed-forward SplatVoxel transformer enables zero-shot history-aware novel view streaming, without training on dynamic scenes.
Table 3: Novel View Streaming Metrics. We compare reconstruction quality of our model to prior methods. Our model differs from these baselines by being history-aware, in contrast to per-frame or temporal methods, and feed-forward rather than optimization-based. Our model outperforms baselines in reconstruction quality and ranks second in temporal smoothness (flicker), whereas the best-ranking baseline in flicker suffers a steep drop in image quality. We report the average running time of processing a single frame on Neural3DVideo at the resolution of $320\times 240$. GSLRM and Ours are tested on a single H100 card. 3DGS, 3DGStream, and 4DGS are tested on a single A100 GPU due to limitations with the H100 card. While the efficiency of these optimization-based methods can be enhanced, they still do not achieve a frame rate of 1 FPS. The best and second-best results are highlighted for clarity.
Category Method Time (s) 4-view LongVolumetricVideo [75] 2-view Neural3DVideo [26]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Flicker $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Flicker $\downarrow$
Per-frame Optimization 3DGS [22] 9.8 18.588 0.6368 0.4334 61.686 21.02 0.6234 0.4989 43.51
3DGStream [62] 3.8 16.970 0.4405 0.5040 12.436 14.24 0.3542 0.6016 6.554
[…]
History-aware Feed Forward Ours 0.07 (non-keyframe) 25.564 0.8645 0.1242 7.782 27.41 0.8863 0.1040 2.916
0.35 (keyframe)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT 3DGS 3DGStream 4DGS GS-LRM Ours
Figure 4: Streaming Reconstruction. We compare against per-frame methods (3DGS and GS-LRM) and temporal methods (4DGS and 3DGStream) on LongVolumetricVideo [75] with 4 input views and Neural3DVideo [26] with 2 input views. 3DGS and 3DGStream struggle with sparse input views, while 4DGS is susceptible to occlusion artifacts. GS-LRM is not designed to handle temporal sequences and prone to flicker, while our method mitigates flickering and produces realistic geometry especially in challenging occluded areas.
4.2.1 3D Warping
Due to scene motion, directly using Gaussian primitives from past frames results in lagging artifacts and fails to capture newly appearing objects. Therefore, we use 2D motion cues from pre-trained models to warp 3D Gaussian primitives from past frames to the current frame.
Triangulation
Given a Gaussian primitive with an initial position $\mu_{t^{\prime}}$ at a previous frame $t^{\prime}$, we estimate its corresponding 3D position $\mu_{t}$ at the current frame $t$ by performing triangulation on the 2D correspondance obtained by a pretrained 2D point tracking model [10]. In detail, we first compute its 2D projection $\{\mathbf{p}_{t^{\prime}}^{i}\mid i=1...N\}$ across $N$ input views. We then operate 2D point tracking on the projection points to obtain its corresponding 2D pixel position $\mathbf{p}_{t}^{i}$ in the current frame. The 3D position is subsequently recovered by finding the closest point to the $N$ camera rays originating from the 2D projections $\mathbf{p}_{t}^{i}$, which can be formulated as the following least-squares optimization problem:
$\centering\mathbf{x}_{c},\lambda_{i}=\arg\min_{\mathbf{x},\lambda_{i}}\sum_{i}\left\|(\mathbf{o}_{i}+\lambda_{i}\mathbf{d}_{i})-\mathbf{x}\right\|^{2}\@add@centering$ (9)
where $\mathbf{o}_{i}$ and $\mathbf{d}_{i}$ denote the origin and direction of the $i$-th camera ray, respectively, and $\lambda_{i}$ is a scalar representing the depth along the ray. This least-squares problems can be further converted into solving an linear equation as
$\centering\left(\sum_{i}\gamma_{i}(I_{3}-\mathbf{d}_{i}\mathbf{d}_{i}^{T})\right)\mathbf{x}=\sum_{i}\gamma_{i}(I_{3}-\mathbf{d}_{i}\mathbf{d}_{i}^{T})\mathbf{o}_{i}\@add@centering$ (10)
where $I_{3}$ denotes the three-dimensional identity matrix, and $\gamma_{i}$ is a visibility mask that indicates whether the projected 2D point is visible throughout the motion occuring in that view. The visibility mask is obtained through a combination of 2D tracking occlusion checks and a depth-based visibility verification, which involves comparing the rendered depth from Gaussian primitives with the projection depth to determine whether the point is visible in front of the surface.
Embedded Deformation Graph
While the algorithm above can estimate 3D motion for any splat, doing so for all primitives is computationally expensive due to 2D tracking, and lacks robustness when splats are occluded. Instead, we use an embedded deformation graph [60, 42, 24] to introduce a spatial continuity prior during online motion estimation. We begin by selecting a subset of anchor splats that are consistently visible across multiple views, using farthest sampling [46]. We then construct the embedded deformation graph using a K-nearest neighbors (KNN) approach, binding each splat to its $K$-nearest splats and computing blend weights based on distance. To warp splats from the previous frame, we computed embedded deformations on the anchor splats using triangulation, and then propagate motion to all primitives through Linear Blend Skinning. For efficiency, we only maintain primitives from a set of past keyframes, and the contribution of these primitives are smoothly adjusted as older keyframes are replaced by newer ones to ensure seamless temporal transitions. Due to time constraints from querying 2D tracking model, in practice we use two keyframes placed 5 frames apart, each lasting for 10 frames.
4.2.2 Error-aware Temporal Fusion
During online reconstruction, we incorporate both historical Gaussian primitives warped from keyframes and new Gaussian primitives predicted from multi-view observation at the current frame. All of the collected Gaussian primitives are deposited to voxels (Sec. 4.1.2) and then processed by the voxel transformer (Sec. 4.1.3), producing refined, history-aware Gaussian primitives.
To further mitigate artifacts resulting from warping inaccuracies of historical primitives (e.g. warping error, appearance variattion, or new objects), we use an adaptive fusion weighting mechanism during splat fusion based on the per-pixel error between the past and current frame splats when rendered to the input views. In more detail, each historical splat is projected onto the input views to assess whether it produces a higher rendering error and appears in front of the rendered depth of the current frame splats. If a historical splat results in increased error in more than half of the input views, we set its fusion weight to zero to suppress unreliable contributions. Concurrently, current frame splats predicted from the multi-view transformer that increase rendering error are assigned zero weight. To enhance temporal continuity, we retain 50% of the voxel attributes from the past keyframe in static areas.
5 Experiments
In this section, we present evaluations of novel view synthesis on both static and dynamic scenes, along with ablation studies on framework design.
5.1 Feed-forward Novel View Synthesis
We first compare our proposed two-stage feed-forward networks with other feed-forward reconstruction methods on the RealEstate10k [82], following the trainining and testing setup in [5]. Since GS-LRM is not open-source, we re-implement it and report its reproduced performance. Table 2 presents a quantitative comparison based on commonly used novel view synthesis metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [70], and the perceptual metric LPIPS [80]. As shown in Fig. 3, our method captures finer details, such as thin structures (e.g., chairs) and reflections, outperforming the state-of-the-art feed-forward Gaussian splatting method, GS-LRM [79].
Additionally, we conduct quantitative comparison with GS-LRM on the DL3DV dataset [37] (Table 4). Our method outperforms GS-LRM by approximately 2dB in PSNR while maintaining a comparable transformer capacity with similar inference time, and using the same training setup. Qualitative comparisons and training details are provided in the Supplementary Document.
Table 4: DL3DV Reconstruction Metrics. We compare our approach with the state-of-the-art feed-forward Gaussian splatting method GS-LRM [37] and present ablation studies on our model designs.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (ms)
GS-LRM 28.59 0.925 0.063 52.8
[…]
Ours (w/o Coarse-to-fine) 20.62 0.587 0.247 48.6
Ours (w/o Sparse Voxel) 29.69 0.926 0.060 72.0
Ours (3D CNN Downsample) 29.44 0.922 0.061 82.1
Ours (w/o Splat Feature) 29.40 0.924 0.061 49.2
[…]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
GT GS-LRM Ours
Figure 5: Reconstrution Close-ups. We show zoomed-in novel view reconstructions of GS-LRM and our model on dynamic scene datasets. Our model better handles occlusion boundaries with sharper detail.
Figure 6: Temporal Slice Comparison. We extract a horizonal line in the ground truth and generated videos to visualize temporal stability. Applying GS-LRM independently per frame is more prone to flickering, while our method achieves greater stability and is more robust to occlusion.
5.2 Novel View Streaming
We compare our method with baselines on the novel view streaming task on two dynamic scene datasets, Neural3DVideo [26] and LongVolumetricVideo [75]. We use 2 input views and 290 frames for the six sequences in Neural3DVideo, and 4 input views with 1500 frames for the two sequences in LongVolumetricVideo. For quantitative evaluation, we use a temporal metric to measure flicker in the rendered streaming videos in addition to the aforementioned per-frame novel view reconstruction metrics. Temporal flickering is measured using feature-space differences over time. Given the predicted frames $I_{t}$ and ground truth frames $I_{t}^{*}$, we extract deep features using a pretrained VGG network [55] $\phi(\cdot)$:
$\text{Flicker}_{t}=\big{|}D_{t}-D_{t}^{*}\big{|}$ (11)
where $D_{t}=\|\phi(I_{t})-\phi(I_{t-1})\|_{2}$ and $D_{t}^{*}=\|\phi(I_{t}^{*})-\phi(I_{t-1}^{*})\|_{2}$ represent the temporal feature differences for the predicted and ground truth frames at frame $t$, respectively. The final flicker score is computed by averaging $\text{Flicker}_{t}$ over the entire sequence. Lower values indicate smoother temporal consistency.
Compared to our proposed history-aware feed-forward method, other baseline methods can be classified into three categories: per-frame optimization, temporal optimization, and per-frame feed-forward reconstruction. As presented in Table 3, our method achieves the highest per-frame novel view synthesis performance across all three evaluation metrics on both datasets with the second-best flicker scores. Notably, our approach significantly enhances novel view synthesis results on the Neural3DVideo dataset, where only two input views are available, demonstrating its robustness in handling extremely sparse input settings. In comparison to the state-of-the-art per-frame feed-forward method GS-LRM, our approach qualitatively reconstructs geometry and appearance with greater accuracy and finer details (Fig. 4), and we show close-ups of the reconstruction in Fig. 5. Our method reduces flickering through history-aware modeling, visualized with temporal slices (Fig. 6). Although 4D Gaussian Splatting achieves a better flicker metric, the novel view quality is substantially worse compared to ours. Moreover, its time-consuming temporal optimization is impractical for our target application, which focuses on online processing of arbitrarily long sequences.
5.3 Ablation Studies
We show model ablations on DL3DV in Table 4. Naive mean averaging of Gaussian attributes in each voxel significantly degrades results. The absence of a coarse-to-fine setup limites the capacity of the output Splats, and without sparsification, we downsampled the grid to 0.75x its original resolution to avoid exceeding memory capacity for the voxel transformer, both leading to poorer reconstruction quality. We also experiment using a 3D CNN to downsample the voxel grid instead of our proposed coarse-to-fine sprase voxel grid, which is $\sim$1 dB worse than our final result and much slower. We also find splat features helpful, operating on splat parameters directly leads to $\sim$1 dB drop in PSNR.
We performed ablation studies on the Neural3DVideo dataset to evaluate our streaming reconstruction components, as shown in Table 5. Removing historical Gaussians fails to resolve temporal flicker. Disabling 3D warping led to abrupt keyframe changes and increased reconstruction errors. Eliminating error-aware fusion degraded reconstruction quality due to warping artifacts and other errors.
Table 5: Ablation Studies on Temporal Components.
[…]
w/o History 27.17 0.890 0.107 5.91
w/o 3D warping 27.12 0.884 0.107 6.09
w/o Error-aware Fusion 26.73 0.874 0.109 3.01
Full Model 27.41 0.886 0.104 2.92
6 Conclusion
In summary, we introduce a novel approach that leverages information from previous frames to enhance novel view streaming with sparse cameras. Our proposed SplatVoxel framework employs a hybrid representation by modeling scenes with Gaussian primitives and fusing them across frames using sparse voxel grids. This design enables training only on static scenes while supporting zero-shot inference on multi-view video streams via explicit 3D scene flow warping. As a result, our method achieves state-of-the-art performance on multiple multi-view video datasets under sparse camera setups and runs at interactive rates on a single H100 GPU.
While our approach marks an important step toward real-time novel view streaming, it does have limitations. Using pretrained 2D tracking models can be slow and prone to errors in challenging scenarios. Moreover, the current framework still performs a partial 3D reconstruction at each frame rather than incrementally updating previous reconstructions. Future work can explore strategies to leverage historical information without explicit tracking and incrementally refine prior reconstructions, potentially prioritizing regions with higher reconstruction errors. More efficient use of historical reconstructions could also improve the handling of high-resolution novel view streaming. These avenues promise to further enhance the efficiency and quality of history-aware novel view streaming.
Acknowledgement.
We would like to express our gratitude to John Flynn, Kathryn Heal, Lynn Tsai, Srinivas Kaza, and Tooba Imtiaz for their helpful discussions, as well as Clément Godard, Jason Lawrence, Michael Broxton, Peter Hedman, Ricardo Martin-Brualla, and Ryan Overbeck for their valuable comments.
Supplementary Material
Appendix A Implementation Details
Our framework is implemented in JAX [3] and trained on NVIDIA A100 GPUs. We evaluate the speed of both GS-LRM and our method on a single NVIDIA H100 GPU. We evaluate the speed of other optimization-based baselines, including 3DGS, 3DGStream, and 4DGS, on a single NVIDIA A100 GPU due to limitations with the H100 card. While running on an H100 can offer speed improvements, these baselines remain constrained by their inherent computational complexity and fail to achieve 1FPS.
We use the Adam optimizer [23] with an initial learning rate of 4e-4, applying cosine learning rate decay with linear warmup. The warmup period is set to 5000 training steps.
A.1 Network Structure
Our network consists of a Multi-view Transformer backbone and a Sparse Voxel Transformer. The multi-view transformer has 24 transformer layers with 1024 hidden dimensions and 16 attention heads. The sparse voxel transformer uses 6 layers with 128 hidden dimensions and 8 attention heads. To demonstrate the effectiveness of our representation, we use a 12-layer multi-view transformer in Table 4, showing that it can outperform GS-LRM by 2dB in PSNR while maintaining a comparable network capacity and similar inference time.
Each transformer layer [66] consists of a self-attention mechanism and feed-forward networks, with all bias terms removed following GS-LRM [79]. Additionally, following LVSM [20], we incorporate QK-Norm [16] to stabilize training. For efficiency, we utilize the CUDNN-FlashAttention [9] implementation available in Jax.
The shallow multi-layer perceptrons $\mathrm{MLP}_{coarse}$ and $\mathrm{MLP}_{fine}$ share similar structures with the feed-forward network in a transformer layer. It includes two linear layers with a LayerNorm [1] and GeLU [15] activation function,and an additional residual connection.
A.2 Hybrid Splat-Voxel Representation
Our SplatVoxel model is inspired by Lagrangian-Eulerian representations in physical simulation where we combine the gaussian splats (Lagrangian representation) and voxels (Eularian representation) to assist warping and fusion of history information.
We use the offical splatting-based rasterizer from 3D Gaussian Splatting [22] to render images from predicted Gaussian primitives. To ensure a fair comparison with GS-LRM on RealEstate10K[82], we set the spherical harmonics degree to 0 for Table 6. For models trained on the DL3DV dataset [37], the degree is set to 1 to enhance view-dependent effects.
Our voxel grid is aligned with the target camera, and is in the normalized device space (NDC), where the z-axis is linear in disparity. This space is helpful in presenting unbounded, in-the-wild scenes. The full fine volume size is set to $[64,H,W]$, where the $H,W$ corresponds to the rendered image height and width. The coarse volume size is set to $[32,H//8,W//8]$, meaning each coarse-level volume is subdivided into $[2,8,8]$ fine-level volumes.
We initialize the voxels by depositing the Gaussian primitives to the volume with a 3D distance-based kernels used in Material Point Method [58]:
$\mathcal{K}(\mathbf{x})=K\left(\frac{x}{h}\right)K\left(\frac{y}{h}\right)K\left(\frac{z}{h}\right),$ (12)
where $\mathbf{x}=(x,y,z)$ is the offset from the voxel center to the Gaussian primitive center, and $h$ is the grid size used to normalize the offset vector to the range $[-1,1]$. The function $K(x)$ is a one-dimensional cubic kernel function given by:
$K(x)=\frac{1}{6}(3|x|^{3}-6x^{2}+4)$ (13)
Rather than directly generating new attributes, we design the voxel transformer to predict residuals that are added to the initial Gaussian rendering attributes on voxels. Specifically, given the input fine-level voxel attributes obtained through the Splat Fusion procedure $\mathbf{E}_{j}$ which comprise the fused splat feature $\mathbf{F}_{j}$ and rendering parameters $\mathbf{G}_{j}$ throught the Splat Fusion procedure, the transformer obtain the final splat rendering parameters $\mathbf{G}^{\prime}_{j}$ in Eqn. 7 as:
$\Delta\mathbf{G}_{j}$ $=\mathrm{MLP}_{fine}([\mathbf{O}_{j},\mathbf{E}_{j}])$ (14)
$\mathbf{G}^{\prime}_{j}$ $=\mathbf{G}_{j}+\Delta\mathbf{G}_{j}.$
A.3 3D Warping
Triangulation
We use TAPIR [10] as our 2D tracking backbone, which takes query points and video frames as input and produces 2D tracking positions throughout the video with the corresponding occlusion mask, indicating whether a tracked point is visible in a given frame. In addition to the occlusion mask obtained from 2D tracking, the visibility mask $\gamma_{i}$ used in Eqn. 10 for triangulation also depends on depth-based visibility. Depth-based visibility is determined by comparing the projected depth of a given splat with the depth map obtained by rendering all splats to a given view. If the projected depth is smaller, the splat is considered unoccluded. This process utilizes alpha blending of splat depths to derive rendered depths from 3D Gaussian Splatting, a widely adopted technique that has been demonstrated to be effective [7].
Embedded Deformation Graph
Assuming motion continuity between neighboring frames, we first estimate optical flow between each keyframe and its subsequent frame using RAFT [64] to identify regions of significant movement. We select regions within the input views where the optical flow norm exceeds a threshold of 0.2 pixels as areas of significant motion. Each splat is then projected onto the input views to determine whether its projected position falls within these significant motion regions and remains unoccluded based on the depth-based visibility verification aforementioned. If a splat exhibits significant motion in more than half of the views, it is prioritized for sampling as an anchor splat. During experiments, we found that uniformly sampling 512 points from the prioritized splats using the Farthest Point Sampling method  [46] is sufficient for effectively handling moving Gaussians in the keyframes.
The deformation of each splat is computed using Linear Blend Skinning (LBS) based on the deformation of its associated anchor splats. For each Gaussian, we employ a K-nearest neighbors approach with $k=4$ to identify the closest anchor splats. The position displacement $\delta\mathbf{p}\in\mathbb{R}^{3}$ is then determined by blending the deformations of the selected anchor points, denoted as $\mathbf{t}_{i}\in\mathbb{R}^{3}$.
$\delta\mathbf{p}$ $=\sum_{i=1}^{k}w_{i}\mathbf{t}_{i}$ (15)
$w_{i}$ $=\frac{\exp(-\|d_{i}\|^{2}/\sigma^{2})}{\sum_{j=1}^{k}\exp(-\|d_{j}\|^{2}/\sigma^{2})},$
where $d_{i}$ represents the Euclidean distance between the Gaussian primitive and the $i$-th anchor point, while $\sigma$ controls the smoothness of the weighting function, set to the mean distance between anchor points and the querying splats. To ensure that static splats remain unaffected, we set displacement $\delta\mathbf{p}$ to be zero if the distance to the anchors is greater than $\lambda\bar{d}$, where $\bar{d}$ is the mean distance between anchor points and $\lambda$ is a hyperparameter controlling the motion graph’s extent. In practice we set $\lambda$ between 0.2 to 0.5 depending on the scene.
Error-aware Temporal Fusion
Static areas are identified during the Splat-to-voxel transfer process, where historical splats undergoing deformation mark their neighboring voxels as regions of motion.
Appendix B Two-stage training
Figure 7: Comparison of Multi-View and Voxel Transformers. Validation curves of our transformer variations on the DL3DV dataset. Compared to the Multi-view transformer which produces gaussian primitives aligned with the 2D input images, our Sparse Voxel Transformer is able to converge at a better result within very few training steps.
Our training procedure first trains the multi-view transformer portion, and then jointly trains the multi-view transformer with the voxel transformer. As shown in the training curves 7, our proposed voxel transformer efficiently reasons about 3D information, surpassing the performance of only using the multi-view transformer with the same number of steps.
Table 6: Reproduced GS-LRM Performance. Our reproduced GS-LRM achieves improved LPIPS and SSIM scores with a slight decrease in PSNR on the RealEstate10K dataset compared to the officially reported numbers.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
pixelNeRF [78] 20.43 0.589 0.550
GPNR [59] 24.11 0.793 0.255
Du et al. [12] 24.78 0.820 0.213
pixelSplat [5] 26.09 0.863 0.136
MVSplat [6] 26.39 0.869 0.128
GS-LRM (Official) [79] 28.10 0.892 0.114
GS-LRM (Reproduced) 27.78 0.902 0.082
Ours 28.47 0.907 0.078
Table 7: Quantitative evaluation comparing our method with the best-performing baselines, 4DGS [73] and GS-LRM [79], across all sequences in LongVolumetricVideo [75] and Neural3DVideo [26]. The best and second-best results are highlighted for clarity.
4DGS GS-LRM Ours
Dataset Sequence PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Flicker $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Flicker $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Flicker $\downarrow$
LongVolCap Corgi 17.25 0.530 0.423 8.17 26.24 0.882 0.118 9.95 26.48 0.885 0.109 8.76
Bike 20.26 0.742 0.288 4.89 24.74 0.830 0.158 6.97 24.64 0.844 0.140 6.80
[…]
cut_roasted_beef 25.41 0.851 0.167 3.17 21.95 0.855 0.134 5.51 29.15 0.909 0.096 3.26
flame_steak 26.54 0.854 0.159 3.24 22.04 0.866 0.121 5.56 29.10 0.903 0.100 3.31
cook_spinach 24.41 0.794 0.198 3.49 21.93 0.872 0.119 5.78 29.26 0.906 0.103 3.42
flame_salmon_1 19.92 0.709 0.250 2.18 21.70 0.827 0.126 5.65 24.70 0.860 0.102 2.52
sear_steak 23.42 0.804 0.206 2.52 22.23 0.869 0.122 4.92 29.14 0.907 0.098 2.64
Appendix C Dataset
C.1 Static Scene Datasets
RealEstate10K [82]
For the RealEstate10K dataset, we set the training and testing resolution to $256\times 256$. Both GS-LRM and our model are trained for 100K iterations. Specifically, we first train our multi-view transformer backbone for 80K iterations, followed by 20K iterations of joint training with the voxel decoder. Each training batch consists of two input views and six target views with a baseline of one unit length between the input views, following the training setup of GS-LRM. For evaluation, we use the same input and target indices as PixelSplat and GS-LRM.
DL3DV [37]
For the DL3DV dataset, we set the training and testing resolution to $384\times 216$. For training, we randomly select one image in the scene as the target, and randomly select four of the nearest eight cameras to the target as inputs. We scale the scene such that the cameras fit within a unit cube. For evaluation, we use every eighth image as the target set, and for each target we use the nearest four cameras not in the target set as inputs. We average metrics per scene, and then average over all scenes. In our experiments on the DL3DV test set (Table 4), both our model and GS-LRM are trained using 8 GPUs, with a batch size of 2 per GPU, for a total of 200K iterations. We first train the multi-view transformer backbone for 100K iterations, followed by 100K iterations of joint training with the voxel decoder.
For cross-dataset evaluation on dynamic scene datasets, we train both GS-LRM and our model on DL3DV using a batch size of 128, for a total of 300K iterations. The training follows a two-stage process: we first train the multi-view transformer backbone for 200K iterations, then jointly train the voxel decoder for 100K iterations.
C.2 Dynamic Scene Datasets
Neural3DVideo [26]
We use six sequences from the dataset for evaluation. The center target camera is selected as the evaluation view, while the two second-closest cameras on either side are used as input cameras. We use resolution $320\times 240$ downsampling from their original resolution.
LongVolumetricVideo [75]
For evaluation, we use the Corgi and Bike sequences and select a 4-input-camera setup that covers one target view. We failed to find enough baseline coverage for other sequences in LongVolCap using only 4 input cameras. For both the Corgi and Bike sequences, we use 1500 frames for evaluation.
We use resolution $384\times 216$ for Corgi and $256\times 256$ for the Bike sequences, downsampling from their original resolution.
Appendix D Baselines
3DGS
We use the original 3D Gaussian Splatting implementation [22], and reduce the number of training iterations from original 30,000 to 1,000 for Neural3DVideo and 1,500 for LongVolumetricVideo. This modification not only accelerates training but also significantly enhances novel view synthesis results (e.g. around 4 PSNR improvement). Given the sparsity of the input data, this adjustment is crucial for preventing overfitting. All other parameters are kept consistent with the original implementation.
3DGStream
We use the open-source 3DGStream [62] implementation to process the multi-view videos. We first use the original 3DGS to optimize on the first frame for 1,500 iterations, warmup the NTC cache, and then optimize on the next frame using the previous 3DGS for 150 iterations, following the original implementation.
4DGS
We use the open-source 4DGS [73] implementation to process the multi-view videos. Following their implementaion, we first optimized the first frame for 3,000 iterations, and then optimize on the full multi-view video sequence for 14,000 iterations.
GS-LRM
We enhance GS-LRM’s speed by replacing the CNN deconvolution with an MLP unpatchify layer as proposed in [20], and implementing it in JAX. We maintain the same network parameters and training loss weights as in the original paper. Additionally, we use the LPIPS loss employed in our model, achieving improved LPIPS and SSIM scores, with a slight decrease in PSNR, as detailed in Table 2. We adopt the same color modeling approach as the official paper for RealEstate10K, setting the spherical harmonics degree to 0. For the GS-LRM baseline trained on DL3DV, we set the degree to 1 to ensure a fair comparison with our method. Additionally, we found that this modification enhances GS-LRM’s performance on DL3DV.
Appendix E Additional Results
Novel View Streaming
Table 7 presents detailed metrics for our results alongside the two best-performing baselines, 4DGS and GS-LRM, evaluated across all eight sequences.
Feed-forward Novel View Synthesis
We also provide additional examples comparing GS-LRM with our method on both DL3DV and RealEstate10K, as shown in Fig. 8 and Fig. 9.
Figure 8: DL3DV Static Scene Reconstruction. Additional qualitative examples of GS-LRM and our method trained on DL3DV.
Figure 9: Re10K Static scene reconstruction. Additional qualitative examples of GS-LRM trained on RealEstate10k [82].